pam is 14 and dad is 37. In how many years will pam's dad be twice as old as she will be?
Answer by mananth(16946)

pam is 14 and dad is 37. In how many years will pam's dad be twice as old as she will be?
let the number of ears be x
37+x= 2*(x+14)
37+x=2x +28
37-28= 2x-x
9=x

After 9 years